Heritage Gardens provides complete garden design and landscaping services. The company uses a job-order costing system to track the costs of its landscaping projects. The table below provides data concerning the three landscaping projects that were in progress during May. There was no work in process at the beginning of May.   

Project
Williams Chandler Nguyen
Designer-hours 300 180 220
Direct materials $ 5,800 $ 2,800 $ 4,600
Direct labor $ 3,400 $ 2,000 $ 2,500


Actual overhead costs were $33,000 for May. Overhead costs are applied to projects on the basis of designer-hours because most of the overhead is related to the costs of the garden design studio. The predetermined overhead rate is $55 per designer-hour. The Williams and Chandler projects were completed in May; the Nguyen project was not completed by the end of the month. No other jobs were in process during May.     

Required:

1. Compute the amount of overhead cost that would have been applied to each project during May.


2. Determine the cost of goods manufactured for May.
Answer #1

1) Overhead applied

Williams Chandler Nguyen Total
Total overhead applied 300*55 = 16500 180*55 = 9900 220*55 = 12100 38500

2) Cost of goods manufactured

Williams Chandler
Direct material 5800 2800
Direct labor 3400 2000
Overhead 16500 9900
Total 25700 14700

Cost of goods manufactured = 25700+14700 = 40400